A tennis ball of mass 0.07 kg is moving towards a racket at a speed of 20 m/s . After collision, the ball moves at a speed of 30 m/s in the opposite direction, as shown below. What is the magnitude of the force acting on the ball if the time of contact between the ball and the racket is 0.01 s?


1
Expert's answer
2021-02-25T11:23:36-0500

The magnitude of force

"F=\frac{\Delta p}{\Delta t}=\frac{m|{\bf v_f-v_i}|}{\Delta t}"

"=\frac{0.07\:\rm kg(30-(-20))\rm m/s}{0.01\:\rm s}=350\:\rm N"